1. (30 points) Bedrock Company has $70 million in debt and $30 million in equity. The debt matures in 1 year and has a 10% interest rate, so the company is promising to pay back $77 million to its debtholders 1 year from now. The company is considering two possible investments, each of which will require an upfront cost of $100 million. Each investment will last for 1 year, and the payoff from each investment depends on the strength of the overall economy. There is a 50% chance that the economy will be weak and a 50% chance it will be strong. Here are the expected payoffs (all dollars are in millions) from the two investments: Payoff in 1 Year If the Economy Is Weak $90.00 Payoff in 1 Year If the Economy Is Strong $130.00 170.00 Expected Payoff Investment L Investment H $110.00 110.00 50.00 Note that the two projects have tlhe same expected payoff, but Project H has higher risk. The debtholders always get paid first and the stockholders receive any money that is available after the debtholders have been paid.
a. (8) Assume that the company selects Investment L. What is the expected payoff to the firms debtholders? What is the expected payoff to the firms stockholders? b. (8) Assume that the company selects Investment H. What is the expected payoff to the firms debtholders? What is the expected payoff to the firms stockholders? (4) Would the debtholders prefer that the companys managers select Project L or Project H? Briefly explain your reason. c. (5) Explain why the companys managers, acting on behalf of the stockholders, might select Project H, even though it has greater risk. d.
(5) What actions can debtholders take to protect their interests? e.
Answer #1

Part a)

The expected payoff to the firm's debtholders and stockholders is calculated as below:

Expected Payoff to Debtholders = Current Value of Debt*(1+Interest Rate) = 70*(1+10%) = $77 million

Expected Payoff to Stockholders = (Payoff in Year 1 if Economy is Weak - Expected Payoff to Debtholders)*Probability of Weak Economy + (Payoff in Year 1 if Economy is Stock - Expected Payoff to Debtholders)*Probability of Strong Economy = (90 - 77)*50% + (130 - 77)*50% = $33 million

_____

Part b)

The expected payoff to the firm's debtholders and stockholders is calculated as below:

Expected Payoff to Debtholders = Probability of Weak Economy*Payoff in Year 1 if Economy is Weak + Probability of Strong Economy*Value of Debt*(1+Interest Rate) = 50%*50 + 50%*70*(1+10%) = $63.5 million

Expected Payoff to Stockholders = (Payoff in Year 1 if Economy is Weak - Expected Payoff to Debtholders)*Probability of Weak Economy + (Payoff in Year 1 if Economy is Stock - Expected Payoff to Debtholders)*Probability of Strong Economy = (50 - 50)*50% + (170 - 77)*50% = $46.50 million

_____

Notes:

If the economy is weak, the entire amount of $50 million will be paid to debtholders. Therefore, stockholders will not get anything in case of a weak economy. However, if the economy is strong, the stockholders will receive the amount ($93 million) that will be left after the debtholders have been paid off.

_____

Part c)

Based on the calculations performed in Part a and Part b, it can be clearly seen that debtholders are better off with investment L. In other words, debtholders would prefer investment L because it is less risky and provides them with a higher return/payoff.

_____

Part d)

Project H provides better return/payoff/cash flow (which is $46.50 million) to shareholders when compared with Project L (which is $33 million). It is because of this reason that managers (acting on behalf of shareholders) may prefer Project H over Project L even when Project H carries higher risk.

_____

Part e)

Restrictive covenants can be included in the bond/loan/debt agreements by the debtholders. Such covenants can restrict managers from procuring more debt, offering higher interest rates to attract more debt, paying dividends to shareholders, undertaking risky investments/projects (such as investment H in this case) and so on.